Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability doubt : Need More explanation

Expert replies
by bitsho » Sat Jul 24, 2010 1:05 pm
Ques : A person has 3 children with atleast one boy . Find the probability of having atleast 2 boys among children (Source : Arun Sharma , Quant Apt for CAT, Problem:14.5)

Solution:

The event is occuring under the following situations :

event 1 :2nd is a boy & 3rd is a girl OR
event 2 : 2nd is a girl and 3rd is a boy OR
event 3 :2nd is a boy & 3rd is a boy

i.e. required prob = 1/2*1/2+1/2*1/2*+1/2*1/2 = 3/4

However,I am unable to understand why we are calculating the prob for event 1 (2nd is a boy & 3rd is a girl ) & event 2 (2nd is a girl and 3rd is a boy) . In my opinion , both represent the same . since the ques is asking about prob of atleast 2 boys , how does the order of girl-boy came in to pic . As per my understanding , prob should be

event 1 :2nd is a boy & 3rd is a girl OR
event 2 : 2nd is a boy & 3rd is a boy

i.e. required prob = 1/2*1/2+1/2*1/2 = 1/2

Please throw some light where i am missing the point .

Any help is appreciated
Last edited by bitsho on Sat Jul 24, 2010 2:21 pm, edited 1 time in total.
Join the discussion
Source: — Problem Solving |

by akdayal » Sat Jul 24, 2010 1:57 pm
Ques : A person has 3 children with atleast one boy . Find the probability of having atleast 2 boys among children (Source : Arun Sharma , Quant Apt for CAT, Problem:14.5)
I also have doubt on answer.
According to question
Total number of ways/events: 1B2G, 2B1G, 3B0G ( As it mentioned in question that person has 3 children with atleast 1 B)
Hence No of fav events: 2B1G, 3B0G (atleast 2B among children)

Hence req. prob = 2/3

I do not think order will matter in this case. If it is can anyone tell me why?
Join the discussion

by GMATGuruNY » Sat Jul 24, 2010 3:02 pm
bitsho wrote:Ques : A person has 3 children with atleast one boy . Find the probability of having atleast 2 boys among children (Source : Arun Sharma , Quant Apt for CAT, Problem:14.5)

Solution:

The event is occuring under the following situations :

event 1 :2nd is a boy & 3rd is a girl OR
event 2 : 2nd is a girl and 3rd is a boy OR
event 3 :2nd is a boy & 3rd is a boy

i.e. required prob = 1/2*1/2+1/2*1/2*+1/2*1/2 = 3/4

However,I am unable to understand why we are calculating the prob for event 1 (2nd is a boy & 3rd is a girl ) & event 2 (2nd is a girl and 3rd is a boy) . In my opinion , both represent the same . since the ques is asking about prob of atleast 2 boys , how does the order of girl-boy came in to pic . As per my understanding , prob should be

event 1 :2nd is a boy & 3rd is a girl OR
event 2 : 2nd is a boy & 3rd is a boy

i.e. required prob = 1/2*1/2+1/2*1/2 = 1/2

Please throw some light where i am missing the point .

Any help is appreciated
Remember:

P(good outcome) = 1 - P(bad outcome)

It's given that we have at least one boy. The question is really asking:

What is the probability that at least 1 of the other 2 children is a boy?

The only bad outcome is if the two other children are girls: 1/2 * 1/2 = 1/4.
So P(at least one additional boy) = 1 - 1/4 = 3/4.

Also, remember that:

P(A and B) = P(A) * P(B)

To determine the probability of events happening together, we have to determine the probability of each separate event and multiply -- and we have to account for all the different ways those separate events could occur.

So P(at least 1B) = P(B) * P(G) + P(G) * P(B) + P(B) * P(B)

Of course, it's much simpler to use the method I described above.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by bitsho » Sat Jul 24, 2010 3:33 pm
Thanks Mitch ! thanks a tonne . so the take away from this problem is

we have to account for all the different ways those separate events could occur.

Thanks again for clarifying this point
Join the discussion

by Stuart@KaplanGMAT » Sat Jul 24, 2010 8:00 pm
bitsho wrote:Ques : A person has 3 children with atleast one boy . Find the probability of having atleast 2 boys among children (Source : Arun Sharma , Quant Apt for CAT, Problem:14.5)
Hi,

where does it say that there's an equal chance that any child is a boy or girl? All the solutions assume that there's a 50/50 chance, but it doesn't say so explicitly.

Without that information, the question is unanswerable.
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by bitsho » Sun Jul 25, 2010 2:40 am
Thanks Stuart ! you have pointed out a basic flaw in question :) I do agree with your point
Join the discussion